OpenStudy (anonymous):

Let a_n = (10^n-1)/9. Define d_n to be the greatest common divisor of a_n and a_n+1. What is the maximum possible value that d_n can take on?

OpenStudy (thomas5267):

That looks very familiar. I think I saw this on this website today.

OpenStudy (anonymous):

gr8, can you please copy and paste that will hel

OpenStudy (thomas5267):

I have no idea where the thread went.

OpenStudy (thomas5267):

\[ a_1=1\\ a_2=11\\ a_3=111 \]

OpenStudy (thomas5267):

It seems that the \(d_n=1\) but I can't prove it.

OpenStudy (anonymous):

if that's the right ans

OpenStudy (thomas5267):

@ganeshie8

OpenStudy (anonymous):

ok.. got it it right thanks

OpenStudy (thomas5267):

AH! \[ 10^{n+1}-1-10(10^n-1)=r\\ 10^{n+1}-1-10^{n+1}+10=r\\ r=9 \] The greatest common divisor of \(10^{n+1}-1\) and \(10^n-1\) is 9. Hence, the greatest common divisor of \(\dfrac{1}{9}\left(10^{n+1}-1\right)\) and \(\dfrac{1}{9}\left(10^n-1\right)\) is 1.

ganeshie8 (ganeshie8):

Nice :)